OpenStudy (anonymous):

Which does not show a direct variation between x and y? A. y=4/x B. y=0.5x C. y=7x D. y=x/9

OpenStudy (anonymous):

@SkaterBoyShawn

OpenStudy (anonymous):

d

OpenStudy (anonymous):

okay

OpenStudy (ajanijones):

Yep it should be D,

OpenStudy (anonymous):

Yay thanks @SkaterBoyShawn and @ajanijones

OpenStudy (anonymous):

no big deal

OpenStudy (anonymous):

Do you think you guys could handle a couple of more or just one more?

OpenStudy (anonymous):

@SkaterBoyShawn and @ajanijones

OpenStudy (ajanijones):

Sure.

OpenStudy (ajanijones):

1 second I'll be back to help you out.

OpenStudy (ajanijones):

Alright I am back.

OpenStudy (anonymous):

Which equations show a direct variation between x and y when solved for y? Choose all answers that are correct. A. -10+ y= 3x - 10 B. y= - 1/8x C. y+9=9x D. y= x-1

OpenStudy (anonymous):

c